OpenStudy (anonymous):

A train traveled 224 miles in 4 hours. At this rate, how far could it travel in 6 hours? A. 226 mi B. 230 mi C. 234 mi D. 336 mi

OpenStudy (anonymous):

I need help with figuring this out will become a fan and give medals

OpenStudy (texaschic101):

make a proportion.. 224/4 = x/6 (224 miles to 4 hrs = x miles to 6 hrs) cross multiply (4)(x) = (224)(6) 4x = 1344 x = 1344/4 x = 336 miles

OpenStudy (anonymous):

Oh ok I get it ty

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

you would take 224 divided by 4 times 6 and you would get 336miles

OpenStudy (anonymous):

hi guys